OpenStudy (anonymous):

the longer side of a triangle is twice as long as the shortest side and 2 cm. longer than the third side. of the perimeter of the triangle is 33 cm, what is the length of each side?

OpenStudy (anonymous):

X+Y+Z=33 X=2Y X=Z+2 X+(X/2)+(X-2)=33 2X+(X/2)=35 5X=70 X=14 Y=7 Z=12